# Cosmic shear with Einstein rings

### v1

@@ Abstract
Cosmic shear is a key probe of modern cosmology. Amongst its challenges are shape noise and intrinsic alignments, both due to our ignorance of the unlensed shape of the source galaxies. I argue here that Einstein rings may be used as standard shape to measure the external shear along their line of sight. In the Euclid era, this new observable is expected to be a competitive and complementary probe of the large-scale structure of the Universe.
@@ 1 Introduction and motivation: cosmic shear is difficult
$C_{\ell}^{\varepsilon}=C_{\ell}^{\gamma}=\left(\frac{3}{2}H_{0}^{2}\Omega_{\rm m}\right)^{2}\int_{0}^{\infty}\mathrm{d}\chi\;W^{2}(\chi)\,P_{\rm m}\left[\chi,\frac{\ell+1/2}{\chi}\right].$ (1)
@@
The aforementioned two difficulties would be alleviated if we disposed of standard shapes, i.e. objects whose intrinsic shape could be inferred independently of weak-lensing distortions. Our research programme, inspired from Birrer et al.,[2] aims to demonstrate that strong-lensing Einstein rings could play this role and be used as complementary probes of cosmic shear.
@@ 2 Modelling the weak lensing of strong lensing
with $\boldsymbol{\beta}^{\prime}=\boldsymbol{\mathcal{A}}_{\rm od}\boldsymbol{\mathcal{A}}_{\rm ds}^{-1}\boldsymbol{\beta}$, $\psi_{\rm eff}(\boldsymbol{\theta})=\psi(\boldsymbol{\mathcal{A}}_{\rm od}\boldsymbol{\theta})$, and $\boldsymbol{\mathcal{A}}_{\rm LOS}=\boldsymbol{\mathcal{A}}_{\rm od}\boldsymbol{\mathcal{A}}_{\rm ds}^{-1}\boldsymbol{\mathcal{A}}_{\rm os}$. This new lens equation is minimal in the sense that it trades three line-of-sight distortion matrices for a single one. Note that eq. (3) is formally equivalent to a lens with potential $\psi_{\rm eff}$ plus perturbations $\boldsymbol{\mathcal{A}}_{\rm LOS}$ all located in the main lens plane.
@@ 4 Towards applications in cosmology
Current photometric galaxy surveys rely on two observables: the fluctuations of the galaxy number density $\delta_{\rm g}$ on the one hand, and their apparent shape $\varepsilon$ on the other hand. From those two observables, one can schematically form three two-point correlation functions: $\langle\delta_{\rm g}\times\delta_{\rm g}\rangle$ (galaxy clustering); $\langle\varepsilon\times\varepsilon\rangle$ (cosmic shear); and $\langle\delta_{\rm g}\times\varepsilon\rangle$ (galaxy–galaxy lensing). Cosmological constraints are then extracted from the analysis of this so-called $3\times 2\text{pt}$ correlation scheme.
The Euclid survey is expected to detect $\mathcal{O}(10^{5})$ strong lenses [7] in addition to the $\mathcal{O}(10^{9})$ galaxy sample. This will add a third observable, the LOS shear $\gamma_{\rm LOS}$, to $\delta_{\rm g}$ and $\varepsilon$. From this third observable, another three two-point correlation functions can be formed – $\langle\gamma_{\rm LOS}\times\gamma_{\rm LOS}\rangle$, $\langle\gamma_{\rm LOS}\times\delta_{\rm g}\rangle$ and $\langle\gamma_{\rm LOS}\times\varepsilon\rangle$ – thereby leading to a $6\times 2\text{pt}$ correlation scheme. The additional source of information coming from $\gamma_{\rm LOS}$ is expected to lead to tighter constraints on the cosmological parameters, while mitigating systematic biases due to intrinsic alignments.
@@ Acknowledgments
Huge thanks to the organisers of this meeting, but also to its participants, for making it so wonderful. I acknowledge support from the French Agence Nationale de la Recherche through the ELROND project (ANR-23-CE31-0002).

### v2

@@ Abstract
Cosmic shear is a key probe of modern cosmology. Amongst its challenges are shape noise and intrinsic alignments, both due to our ignorance of the unlensed shape of the source galaxies. I argue here that Einstein rings may be used as standard shapes to measure the external shear along their line of sight. In the Euclid era, this new observable is expected to be a competitive and complementary probe of the large-scale structure of the Universe.
@@ 1 Introduction and motivation: cosmic shear is difficult
$C_{\ell}^{\varepsilon}=C_{\ell}^{\gamma}=\left(\frac{3}{2}H_{0}^{2}\Omega_{\rm m}\right)^{2}\int_{0}^{\infty}\mathrm{d}\chi\;W^{2}(\chi)\,P_{\rm m}\left(\chi,\frac{\ell+1/2}{\chi}\right).$ (1)
@@
The aforementioned two difficulties would be alleviated if we disposed of standard shapes, i.e. objects whose intrinsic shape could be inferred independently of weak-lensing distortions. Our research programme, inspired by Birrer et al.,[2] aims to demonstrate that strong-lensing Einstein rings could play this role and be used as complementary probes of cosmic shear.
@@ 2 Modelling the weak lensing of strong lensing
with $\boldsymbol{\beta}^{\prime}=\boldsymbol{\mathcal{A}}_{\rm od}\boldsymbol{\mathcal{A}}_{\rm ds}^{-1}\boldsymbol{\beta}$, $\psi_{\rm eff}(\boldsymbol{\theta})=\psi(\boldsymbol{\mathcal{A}}_{\rm od}\boldsymbol{\theta})$, and $\boldsymbol{\mathcal{A}}_{\rm LOS}=\boldsymbol{\mathcal{A}}_{\rm od}\boldsymbol{\mathcal{A}}_{\rm ds}^{-1}\boldsymbol{\mathcal{A}}_{\rm os}$. This new lens equation is minimal in the sense that it trades three line-of-sight distortion matrices for a single one. Note that eq. (3) is formally equivalent to a main lens with potential $\psi_{\rm eff}$ plus perturbations $\boldsymbol{\mathcal{A}}_{\rm LOS}$ all located in the main lens plane.
@@ 4 Towards applications in cosmology
Current photometric galaxy surveys rely on two observables: the fluctuations of the galaxy number density $\delta_{\rm g}$ on the one hand, and their apparent shapes $\varepsilon$ on the other hand. From those two observables, one can schematically form three two-point correlation functions: $\langle\delta_{\rm g}\times\delta_{\rm g}\rangle$ (galaxy clustering); $\langle\varepsilon\times\varepsilon\rangle$ (cosmic shear); and $\langle\delta_{\rm g}\times\varepsilon\rangle$ (galaxy–galaxy lensing). Cosmological constraints are then extracted from the analysis of this so-called $3\times 2\text{pt}$ correlation scheme.
The Euclid survey is expected to detect $\mathcal{O}(10^{5})$ strong lenses in addition to the $\mathcal{O}(10^{9})$ galaxy sample.[7] This will add a third observable, the LOS shear $\gamma_{\rm LOS}$, to $\delta_{\rm g}$ and $\varepsilon$. From this third observable, another three two-point correlation functions can be formed – $\langle\gamma_{\rm LOS}\times\gamma_{\rm LOS}\rangle$, $\langle\gamma_{\rm LOS}\times\delta_{\rm g}\rangle$ and $\langle\gamma_{\rm LOS}\times\varepsilon\rangle$ – thereby leading to a $6\times 2\text{pt}$ correlation scheme. The additional source of information coming from $\gamma_{\rm LOS}$ is expected to lead to tighter constraints on the cosmological parameters, while mitigating systematic biases due to intrinsic alignments.
@@ Acknowledgments
I would like to thank the organisers of the 2024 58th Rencontres de Moriond, but also its participants, for making this meeting so wonderful. I acknowledge support from the French Agence Nationale de la Recherche through the ELROND project (ANR-23-CE31-0002).